# DiagramQG: A Dataset for Generating Concept-Focused Questions from Diagrams

### v1

@@ 5.1 Implementation Details
The experimental framework uses T5 Base and Large architectures along with the CLIP (ViT-B/32) model. Optimization is performed with the AdamW optimizer, applying a learning rate of 1e-5 for CLIP and 5e-5 for other components, running through 20 epochs of fine-tuning. Key parameters include a maximum input sequence length of 256, an output sequence length of 32, and a batch size of 32, with experiments conducted on one NVIDIA Tesla A800 80G GPU.
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 Bert-Score METEOR CIDEr ROUGE-L DEHN
In-Context Learning
InternVL2 (8B) [3] 25.57 14.48 8.76 4.54 75.18 12.78 0.25 22.12 0.37
InternVL2 (26B) [3] 27.84 16.54 10.76 7.64 78.72 15.56 0.37 24.12 0.45
InternVL2 (40B) [3] 28.40 17.78 11.85 8.63 79.72 16.91 0.39 25.35 0.46
MiniCPM-V2(8B) [48] 24.25 12.51 7.86 4.89 74.64 11.49 0.28 21.96 0.45
DeepSeek-VL(7B) [21] 23.96 14.14 8.53 5.51 76.12 14.96 0.30 23.42 0.47
Qwen2-VL(2B) [39] 26.87 15.53 9.35 5.59 75.16 12.76 0.29 23.21 0.34
Qwen2-VL(7B) [39] 28.64 17.71 11.51 8.59 76.68 15.18 0.45 24.79 0.36
Qwen2-VL(72B) [39] 35.82 25.17 18.94 15.21 82.19 19.84 0.71 31.76 0.42
GLM4-V pro 31.89 20.05 13.87 10.68 81.16 16.52 0.62 27.41 0.37
Claude-3.5-Sonnet (T) 22.76 13.01 8.05 5.16 78.77 13.26 0.45 20.51 0.02
GPT-4o (T) 19.31 10.22 6.17 3.65 80.54 11.02 0.36 20.7 0.02
Claude-3.5-Sonnet 41.86 35.15 30.14 26.20 83.54 30.39 1.78 39.67 0.20
GPT-4o 43.72 37.22 32.26 28.13 84.16 31.50 1.91 43.39 0.23
Fine-tuned
K-VQG [34] 22.37 16.97 13.51 11.05 73.88 27.48 0.45 28.55 0.25
Patch-TRM [22] 23.02 17.63 14.08 11.67 74.68 28.74 0.47 30.50 0.28
ConVQG (BLIP) [25] 31.21 23.42 18.47 15.06 74.71 29.89 0.56 31.24 0.43
KC-VQG (BLIP) [18] 33.13 26.56 21.75 17.52 79.84 31.59 1.03 34.26 0.45
KC-VQG (Qwen2-VL 2B) [18] 36.11 28.74 23.59 19.72 83.54 33.93 1.32 37.54 0.54
Ours
HKI-DQG(BLIP, T5-Base) 47.98 40.25 35.31 31.87 84.27 41.68 2.91 46.31 0.29
HKI-DQG(Qwen2-VL 2B, T5-Base) 58.76 52.64 48.52 46.94 88.74 52.44 4.15 58.21 0.51
HKI-DQG(Qwen2-VL 7B, T5-Base) 60.62 54.68 50.54 47.48 89.52 53.09 4.19 59.97 0.57
HKI-DQG(BLIP, T5-Large) 50.63 43.61 38.94 35.69 85.44 43.84 3.11 49.75 0.31
HKI-DQG(Qwen2-VL 2B, T5-Large) 60.03 54.14 50.07 47.05 89.28 53.21 4.21 59.62 0.54
HKI-DQG(Qwen2-VL 7B, T5-Large) 61.63 55.62 51.46 48.41 89.81 54.26 4.32 60.69 0.59
Table 2: Results on DiagramQG, where GPT-4o (T) and Claude-3.5-Sonnet (T) are meant to generate questions using only text constraints.
@@ 5.2 Baselines
VQG approaches can be classified into two main categories: Fine-tuned and In-Context Learning methods. Fine-tuned methods use models like BERT [6], T5 [33], or GPT-2 [31] for question generation, with examples including K-VQG [34], Patch-TRM [22], ConVQG [25], and KC-VQG [18]. In-Context Learning methods utilize large vision-language models to extract relevant information from diagrams and generate questions using three reference questions. This includes open-source models like Qwen2-VL [39], MiniCPM-V [48], DeepSeek-VL [21], and InternVL2 [3], as well as closed-source models such as GLM4-V pro, Claude-3.5 Sonnect, and GPT-4o. Models such as BLIP2 [13], Llava 1.5 [17], and CogVLM [42] do not perform satisfactorily on the DiagramQG dataset, likely caused by limited exposure to diagram datasets.
@@
Figure 8: Four different cases in DiagramQG dataset.
@@ 5.3 Evaluation metrics
We evaluate our model using established language generation metrics, including BLEU [28], Bert Score, METEOR [5], CIDEr [37], and ROUGE [15]. We evaluate model performance using the pycocoevalcap package to measure alignment between generated and ground truth questions. In addition to traditional question metrics, we propose the Diagram Element Hit Number (DEHN) to measure the utilization of diagram elements in the generated questions. Due to the uneven distribution of elements across diagrams, with some containing over 30 elements, a single question cannot cover all elements. Thus, we define DEHN as the count of diagram elements present in the generated question, providing a complementary metric to assess the model’s ability to incorporate relevant diagram information. The DEHN is calculated as follows: $\text{DEHN}=\frac{1}{|Q|}\sum_{q\in Q}\sum_{i=1}\mathbb{1}(e_{i}\in q)$ where $|Q|$ represents the total number of generated questions, $q$ is a single question from the set of generated questions $Q$, and $\mathbb{1}(e_{i}\in q)$ is an indicator function that equals 1 if the diagram element $e_{i}$ is present in the question $q$ and 0 otherwise. This metric quantitatively measures how well the model incorporates relevant diagram elements into the generated questions, complementing traditional metrics.
@@ 5.4 Results
We evaluate various models on our DiagramQG dataset, as shown in Table 2. Among the in-context learning methods that utilize Vision-Language Models (VLMs), larger models generally perform better, indicating their abundant knowledge across different domains. Interestingly, Qwen2-VL 2B achieves impressive results despite having fewer parameters, making it a suitable base model for further experiments. Tests conducted with GPT-4o and Claude-3.5 Sonnet using only text inputs result in both lower DEHN scores and reduced generation performance, highlighting the importance of incorporating diagrams in the question generation process. Among the fine-tuned methods, KVQG and Patch-TRM do not leverage VLMs for additional knowledge, while ConVQG and KC-VQG do.
The latter models demonstrate superior performance, with KC-VQG showing improved results as VLM parameters scale up. HKI-DQG framework achieves the best performance across all metrics, regardless of whether it is based on T5-base or T5-large, particularly when using the larger-scale Qwen2-VL model. These impressive results validate HKI-DQG as an effective baseline for future research by demonstrating the value of incorporating subject knowledge in the DQG task. We further analyze the performance of HKI-DQG across different concepts, as illustrated in Figure 7. The results exhibit a long-tail distribution, where concepts with more training samples generally perform better.
@@ 5.5 Results on Other Dataset
We evaluate HKI-DQG’s generalizability on two VQG datasets of natural images, namely VQG-COCO and K-VQG. Our proposed HKI-DQG consistently outperforms previous state-of-the-art methods on both VQG-COCO and K-VQG datasets, as shown in Table 3.
Method BLEU-1/BLEU-4 METEOR CIDEr
IM-VQG [11] -/11.44 -/17.07 –/0.26
MC-BMN [29] 40.7/- 22.6/- 0.50/-
K-VQG [34] -/18.84 -/22.79 -/1.31
ConVQG [25] 50.2/20.01 26.4/22.66 0.56/1.53
[…]
Table 3: Evaluation results on VQG-COCO (left of /) and K-VQG (right of /) datasets.
Setting BLEU-4 METEOR CIDEr DERN
n=1,m=4 40.69 45.93 3.31 0.35
n=2,m=4 43.76 50.68 3.64 0.43
n=4,m=4 46.12 51.95 4.02 0.53
n=3,m=0 42.82 49.14 3.72 0.42
n=3,m=1 43.63 49.81 3.82 0.46
n=3,m=2 44.59 50.26 3.88 0.49
n=3,m=3 45.65 51.70 4.05 0.51
n=3,m=5 46.69 52.28 4.12 0.52
n=3,m=4 46.94 52.44 4.15 0.51
@@ 5.6 Parameter Sensitive Study
Our experiments involve two manually set parameters: dividing the diagram into an $n$-layer pyramid of patches and selecting $m$ sentences of background knowledge for question generation. Table 4 shows the results with $n$ ranging from $1$ to $4$, with the model performance peaking at $n=3$, suggesting that overly fine diagram division leads to noisy knowledge extraction by VLM, affecting model performance. For the second parameter, Table 4 shows results for m from $1$ to $5$, peaking at $m=4$. While excessive knowledge affects the performance of HKI-DQG, the minor impact demonstrates the framework’s resilience to redundant inputs.
@@ 5.7 Case Study
The evaluation of the HKI-DQG framework on the DiagramQG dataset (Figure 8) reveals good capability in question generation, but also highlights areas for improvement. We summarize three classes for the poor performance cases: Text Constraint Lost, Diagram Interpretation Bias, and Concept Ambiguity. The first case (Text Constraint Lost) exposes a limitation in text constraint, where HKI-DQG generates questions based primarily on visual features while ignoring the target text. In the second case (Diagram Interpretation Bias), HKI-DQG generates a valid question but fails to capture the complex ecological interactions, only employing the simple relationship between bivalves and sea ducks. The third case (Concept Ambiguity) demonstrates poor semantic alignment between the generated question and the specific concept. These findings suggest that while HKI-DQG shows basic functionality, improvements in text constraint adherence, diagram understanding, and conceptual semantic alignment are needed. For a more detailed case study analysis, please refer to the supplementary materials.
@@ 6 Conclusion
This study introduces DiagramQG dataset, the first dataset for diagram question generation. We propose Hierarchical Knowledge Integration for Diagram Question Generation (HKI-DQG) framework to generate concept-focused questions, supporting interactive and personalized learning [26].
@@ B. More details on baselines
GLM4-V is a large-scale multimodal language model developed by ZhiYuan, featuring outstanding visual understanding and language generation capabilities. It employs a unified pre-training framework, capable of handling multiple modalities such as text, images, and audio. GLM4-V demonstrates strong performance in tasks like visual question answering, image description, and cross-modal reasoning, and can quickly adapt to new scenarios through few-shot learning. The model supports both Chinese and English, excelling in multilingual understanding and generation.
Claude 3.5 Sonnect is a next-generation multimodal assistant developed by Anthropic, exhibiting exceptional performance in both visual and language understanding. It utilizes an innovative neural network architecture that allows for deep comprehension of image content and complex reasoning. The model has strict controls in terms of safety and ethics, capable of identifying and filtering inappropriate content. A notable feature is its strong contextual understanding and coherent conversational abilities.
GPT-4 is the latest large language model developed by OpenAI, possessing powerful multimodal understanding and generation capabilities. It can process image and text inputs and perform complex reasoning and knowledge integration. This model showcases remarkable few-shot learning abilities, quickly mastering new tasks with only a few examples. GPT-4 also excels in creativity, generating high-quality text, code, and other creative content.
@@
BLIP2 [13], Llava 1.5 [17], and CogVLM [42], despite their strong performance in general vision-language tasks, showed limitations in diagram-specific question generation, possibly due to insufficient training on diagram-focused datasets.
@@ D. More experimental details on other VQG datasets
We evaluated our baseline model’s generalizability in visual question generation on two datasets of natural images, namely VQG-COCO and K-VQG.
@@
K-VQG is a knowledge-aware VQG dataset that uniquely combines visual content with structured knowledge triplets. It is a large-scale, human-annotated dataset containing around 13,000 images, 16,000 question-answer pairs, and 6,000 knowledge triplets. Each sample consists of an image, a question, an answer, and a knowledge triplet. K-VQG allows the use of both answers or knowledge triplets as text constraints.
Our proposed HKI-DQG consistently achieves substantial improvements over previous state-of-the-art methods across both VQG-COCO and K-VQG datasets, as shown in Table 6 and 7. These results demonstrate the strong generalization capability of our approach and prove that using a more advanced vision-language model as a foundation can significantly enhance performance in visual question generation tasks.

### v3

@@ 5.1 Implementation Details
The experimental framework uses T5 Base and Large architectures along with the CLIP (ViT-B/32) model. Optimization is performed with the AdamW optimizer, applying a learning rate of 1e-5 for CLIP and 5e-5 for other components, running through 20 epochs of fine-tuning. Key parameters include a maximum input sequence length of 256, an output sequence length of 32, and a batch size of 32, with experiments conducted on one NVIDIA A800 80G GPU. The DiagramQG dataset is split into training, validation, and testing sets with a ratio of 70:5:25, ensuring no overlap of diagrams and questions across these sets. Additionally, each concept is represented in both validation and testing sets, despite some concepts having relatively few samples.
Method BLEU-1 BLEU-2 BLEU-3 BLEU-4 Bert-Score METEOR CIDEr ROUGE-L In-Context Learning InternVL2 (26B) [3] 27.84 16.54 10.76 7.64 78.72 15.56 0.37 24.12 InternVL2 (40B) [3] 28.40 17.78 11.85 8.63 79.72 16.91 0.39 25.35 MiniCPM-V2(8B) [47] 24.25 12.51 7.86 4.89 74.64 11.49 0.28 21.96 DeepSeek-VL(7B) [20] 23.96 14.14 8.53 5.51 76.12 14.96 0.30 23.42 Qwen2-VL(2B) [38] 26.87 15.53 9.35 5.59 75.16 12.76 0.29 23.21 Qwen2-VL(7B) [38] 28.64 17.71 11.51 8.59 76.68 15.18 0.45 24.79 Qwen2-VL(72B) [38] 35.82 25.17 18.94 15.21 82.19 19.84 0.71 31.76 GLM4-V pro 31.89 20.05 13.87 10.68 81.16 16.52 0.62 27.41 Claude-3.5-Sonnet (T) 22.76 13.01 8.05 5.16 78.77 13.26 0.45 20.51 GPT-4o (T) 19.31 10.22 6.17 3.65 80.54 11.02 0.36 20.7 Claude-3.5-Sonnet 41.86 35.15 30.14 26.20 83.54 30.39 1.78 39.67 GPT-4o 43.72 37.22 32.26 28.13 84.16 31.50 1.91 43.39 Fine-tuned K-VQG [33] 22.37 16.97 13.51 11.05 73.88 27.48 0.45 28.55 Patch-TRM [21] 23.02 17.63 14.08 11.67 74.68 28.74 0.47 30.50 ConVQG (BLIP) [24] 31.21 23.42 18.47 15.06 74.71 29.89 0.56 31.24 KC-VQG (BLIP) [17] 33.13 26.56 21.75 17.52 79.84 31.59 1.03 34.26 KC-VQG (Q-2B) [17] 36.11 28.74 23.59 19.72 83.54 33.93 1.32 37.54 Ours HKI-DQG(BLIP, T5-B) 47.98 40.25 35.31 31.87 84.27 41.68 2.91 46.31 HKI-DQG(Q-2B, T5-B) 58.76 52.64 48.52 46.94 88.74 52.44 4.15 58.21 HKI-DQG(Q-7B, T5-B) 60.62 54.68 50.54 47.48 89.52 53.09 4.19 59.97 HKI-DQG(BLIP, T5-L) 50.63 43.61 38.94 35.69 85.44 43.84 3.11 49.75 HKI-DQG(Q-2B, T5-L) 60.03 54.14 50.07 47.05 89.28 53.21 4.21 59.62 HKI-DQG(Q-7B, T5-L) 61.63 55.62 51.46 48.41 89.81 54.26 4.32 60.69
Table 2: Results on DiagramQG, where GPT-4o (T) and Claude-3.5-Sonnet (T) are meant to generate questions using only text constraints, where, Q-2B, Q-7B, T5-B, T5-L mean, Qwen2-VL-2B, Qwen2-VL-7B, T5-Base and T5-Large.
[…]
Table 3: Evaluation results on K-VQG dataset.
@@ 5.2 Baselines
VQG approaches can be classified into two main categories: Fine-tuned and In-Context Learning methods. Fine-tuned methods use models like BERT [6], T5 [32], or GPT-2 [30] for question generation, with examples including IM-VQG [11], K-VQG [33], Patch-TRM [21], ConVQG [24], LV2-Net [19], and KC-VQG [17]. In-Context Learning methods utilize large vision-language models to extract relevant information from diagrams and generate questions using three reference questions. This includes open-source models like Qwen2-VL [38], MiniCPM-V [47], DeepSeek-VL [20], and InternVL2 [3], as well as closed-source models such as GLM4-V pro, Claude-3.5 Sonnect, and GPT-4o. Models such as Llava 1.5 [16], and CogVLM [41] do not perform satisfactorily on the DiagramQG dataset, likely caused by limited exposure to diagram datasets.
@@ 5.3 Evaluation metrics
We evaluate our model using established language generation metrics, including BLEU [27], Bert Score, METEOR [5], CIDEr [36], and ROUGE [14]. We evaluate model performance using the pycocoevalcap package to measure alignment between generated and ground truth questions.
@@ 5.4 Results
We evaluate various models on our DiagramQG dataset, as shown in Table 2. Among the in-context learning methods that utilize Vision-Language Models (VLMs), models with larger parameter counts generally perform better, indicating their abundant knowledge across different subjects. Interestingly, Qwen2-VL 2B achieves impressive and competitive results despite having fewer parameters, making it a suitable base model for further comprehensive experiments. Tests conducted with GPT-4o and Claude-3.5-Sonnet using only text inputs result in reduced generation performance, highlighting the critical importance of incorporating diagrams in the question generation process. Among the fine-tuned methods, KVQG and Patch-TRM do not leverage VLMs for additional knowledge, while ConVQG and KC-VQG do. The latter models demonstrate superior performance, with KC-VQG showing improved results as VLM parameters scale up. HKI-DQG framework achieves the best performance across all quantitative evaluation metrics, regardless of whether it is based on T5-base or T5-large, particularly when using the larger-scale Qwen2-VL model. These impressive results validate HKI-DQG as an effective baseline for future research by demonstrating the value of incorporating subject knowledge in the DQG task. We further analyze the performance of HKI-DQG across different concepts, as illustrated in Figure 7. The results exhibit a long-tail distribution, where concepts with more training samples consistently and significantly perform better.
[…]
Table 4: Evaluation results on VQG-COCO, OK-VQA and A-OKVQA datasets.
@@ 5.5 Results on Other Dataset
We conduct comprehensive evaluations of HKI-DQG’s generalization capabilities across four prominent VQG datasets of natural images: VQG-COCO, OK-VQA, A-OKVQA, and K-VQG, with results presented in Tables 4 and 3. The experimental results demonstrate that HKI-DQG consistently outperforms existing approaches across most evaluation metrics, with notably superior performance in CIDEr and METEOR scores, suggesting enhanced semantic comprehension and contextually coherent question generation.
H.K.E K.S K.F BLEU-4 METEOR CIDEr
$\checkmark$ $\checkmark$ $\checkmark$ 46.94 52.44 4.15
$\checkmark$ $\checkmark$ 44.65 50.07 3.86
$\checkmark$ $\checkmark$ 40.87 48.53 3.27
$\checkmark$ 39.13 46.87 3.13
$\checkmark$ 16.47 28.24 0.53
Table 5: Results on DiagramQG with Qwen2-VL 2B & T5-Base to obtain the results, where H.K.E, K.S and K.F mean HierKnowExtract, KnowSelect and KnowFusionQG.
5.6 Ablation experiment
Given that KnowSelect is inherently based on HierKnowExtract, isolated evaluation of KnowSelect is not feasible. The ablation study results are presented in Table 5. The findings demonstrate that incorporating relevant knowledge substantially improves question generation quality. Notably, KnowSelect shows greater impact than KnowFusionQG, indicating that while HierKnowExtract generates abundant knowledge, KnowSelect effectively filters and retains the most relevant information.
Setting BLEU-4 METEOR CIDEr
n=1,m=4 40.69 45.93 3.31
n=2,m=4 43.76 50.68 3.64
n=4,m=4 46.12 51.95 4.02
n=3,m=1 42.63 48.81 3.72
n=3,m=2 44.59 50.26 3.88
n=3,m=3 45.65 51.70 4.05
n=3,m=5 46.69 52.28 4.12
n=3,m=4 46.94 52.44 4.15
@@ 5.7 Parameter Sensitive Study
Our experiments involve two manually set key parameters: dividing the diagram into an $n$-layer pyramid of patches and selecting $m$ sentences of background knowledge for question generation. Based on our analysis, Table 6 clearly shows the detailed results with $n$ ranging from $1$ to $4$, with the model performance consistently peaking at $n=3$, strongly suggesting that overly fine diagram division leads to noisy knowledge extraction by VLM, significantly affecting model performance. Regarding the second important parameter, Table 6 systematically shows experimental results for m ranging from $1$ to $5$, definitively peaking at $m=4$. While excessive background knowledge notably affects the performance of HKI-DQG, the relatively minor overall impact clearly demonstrates the framework’s impressive resilience to redundant input information.
Figure 8: Comparison between HKI-DQG and GPT-4o using two distinct examples from the DiagramQG dataset.
@@ 5.8 Case Study
Figure 8 illustrates examples from the DiagramQG dataset comparing the HKI-DQG framework and GPT-4o, showcasing HKI-DQG’s superior question generation capabilities. While advanced models like GPT-4o can generate coherent and fluent question texts, those questions often fail to effectively test students’ deeper understanding of concepts. Compared to GPT-4o, HKI-DQG consistently demonstrates significant advantages in generating questions with stronger concept relevance. For a more detailed case study analysis and results, please refer to the supplementary materials.
@@ 6 Conclusion
This study introduces DiagramQG, the first dataset specifically designed for diagram question generation. We propose HKI-DQG, a novel framework that leverages hierarchical knowledge integration to generate concept-focused questions. Experiments show that HKI-DQG framework outperforms existing VQG and popular vision-language models. This work advances intelligent education [25] by enabling automated question generation for personalized learning.
@@ B. More details on baselines
GLM4-V is a large-scale multimodal language model developed by ZhiPu, featuring outstanding visual understanding and language generation capabilities. It employs a unified pre-training framework, capable of handling multiple modalities such as text, images, and audio. GLM4-V demonstrates strong performance in tasks like visual question answering, image description, and cross-modal reasoning, and can quickly adapt to new scenarios through few-shot learning. The model supports both Chinese and English, excelling in multilingual understanding and generation.
Claude 3.5 Sonnet is a next-generation multimodal assistant developed by Anthropic, exhibiting exceptional performance in both visual and language understanding. It utilizes an innovative neural network architecture that allows for deep comprehension of image content and complex reasoning. The model has strict controls in terms of safety and ethics, capable of identifying and filtering inappropriate content. A notable feature is its strong contextual understanding and coherent conversational abilities.
GPT-4o is the latest large language model developed by OpenAI, possessing powerful multimodal understanding and generation capabilities. It can process image and text inputs and perform complex reasoning and knowledge integration. This model showcases remarkable few-shot learning abilities, quickly mastering new tasks with only a few examples. GPT-4o also excels in creativity, generating high-quality text, code, and other creative content.
@@
Llava 1.5 [16], and CogVLM [41], despite their strong performance in general vision-language tasks, showed limitations in diagram-specific question generation, possibly due to insufficient training on diagram-focused datasets.
@@ C. More details on DiagramQG
Figure 12: Three different cases in DiagramQG dataset.
Figure 13: Three different cases in DiagramQG dataset.
@@ D. More experimental details on other VQG datasets
Method BLEU-4 METEOR CIDEr
IM-VQG [11] 36.47 30.25 0.15
KVQG [33] 27.18 55.38 0.13
LV2-Net [19] 29.90 92.17 0.15
[…]
Table 10: Results on the OK-VQA dataset. The results of IM-VQG are reproduced based on the official code.
Method BLEU-4 METEOR CIDEr
IM-VQG [11] 39.30 22.11 0.12
KVQG [33] 30.56 40.97 0.13
LV2-Net [19] 32.11 60.06 0.14
[…]
Table 11: Results on the A-OKVQA dataset. The results of IM-VQG are reproduced based on the official code.
We evaluated our baseline model’s generalizability in visual question generation on four datasets of natural images, namely VQG-COCO, K-VQG, OK-VQA and A-OKVQA datasets.
@@
K-VQG is a knowledge-aware VQG dataset that uniquely combines visual content with structured knowledge triplets. It is a large-scale, human-annotated dataset containing around 13,000 images, 16,000 question-answer pairs, and 6,000 knowledge triplets. Each sample consists of an image, a question, an answer, and a knowledge triplet. K-VQG allows the use of both answers or knowledge triplets as text constraints. OK-VQA represents a groundbreaking dataset in knowledge-based visual question answering, featuring 14,055 open-ended questions with 5 ground-truth answers each. Unlike previous VQA datasets that rely on pre-defined knowledge bases, OK-VQA uniquely requires models to independently retrieve and utilize external knowledge sources like Wikipedia to formulate answers. The dataset has been carefully curated through manual filtering to ensure that every question necessitates external knowledge, with deliberate reduction of frequently occurring answers to minimize dataset bias.
Building upon OK-VQA’s foundation, A-OKVQA emerges as its augmented successor, encompassing approximately 25,000 distinct questions with no overlap in question-image pairs with its predecessor. What sets A-OKVQA apart is its emphasis on comprehensive reasoning skills and diverse knowledge sources, extending beyond simple knowledge base queries. The dataset is specifically designed to challenge models with questions that demand complex commonsense reasoning about visual scenes, requiring the integration of both world knowledge and visual understanding to generate accurate answers.
Our proposed HKI-DQG consistently achieves substantial improvements over previous state-of-the-art methods across both VQG-COCO, K-VQG, OK-VQA and A-OKVQA datasets, as shown in Table 8, 9, 10 and 11. These results demonstrate the strong generalization capability of our approach and prove that using a more advanced vision-language model as a foundation can significantly enhance performance in visual question generation tasks.